# LLM-based feature generation from text for interpretable machine learning

### v1

@@ 3.4.2 Explanation of generated features
Explanation of features with the use of Shapley values was done on two feature subsets for both CORD-19 and M17+, which are the LLM-feature only and BoW + LLM features, where besides the LLM-generated features, the BoW terms were also observed.
In Figure 1, it can be observed that in the case of CORD-19 and prediction of citation rate, the most influential features are basic_medicine, rigor, where higher values indicate higher citation rate.
For the M17+ dataset, the main deciding factor in the evaluation is rigor and grammar, replicablity and novelty. These values can be interpreted as more rigorous, error-free abstracts with high novelty and explicitly mentioned reproducibility abstracts are generally evaluated better.
@@
Figure 1: SHAP plot the GradientBoost model trained on LLM-feature only feature set for CORD-19 (left) and M17+ (right). The X-axis represents the SHAP value, indicating the feature’s effect on the prediction—positive values increase the prediction, while negative values decrease it. Features are ranked by importance on the Y-axis. The color of the points corresponds to the feature values. For binary features like grammar or basic_medicine, blue indicates a value of 0, while red represents a value of 1. For categorical features such as rigor or novelty, blue denotes a low value, purple indicates a medium value, and red signifies a high value.
Similarly to the LLM-feature only subsets, the SHAP analysis was performed on BoW + LLM-features subset to compare the importance of LLM-generated features with the importance of TF-IDF weighted BoW terms. Figure 2 shows that the most important features in both datasets are LLM-generated rather than BoW terms. It can be, however, seen that in the case of CORD-19, there is a larger prevalence of BoW terms in the top 20 most influential variables.
Figure 2: SHAP plot of the GradientBoost model trained on BoW + LLM-features feature set for CORD-19 (left) and M17+ (right). The X-axis represents the SHAP value, indicating the feature’s effect on the prediction—positive values increase the prediction, while negative values decrease it. Features are ranked by importance on the Y-axis. The color of the points corresponds to the feature values. For binary features like grammar or basic_medicine, blue indicates a value of 0, while red represents a value of 1. For categorical features such as rigor or novelty, blue denotes a low value, purple indicates a medium value, and red signifies a high value.
Features marked * were generated by LLM, the remaining features are from the BoW model.
3.5 Action rules
@@ 3.5.1 M17+ Dataset - Bad to Good Evaluation Transition
$r_{9}$ med $\rightarrow$ high 0 $\rightarrow$ 1 bad $\rightarrow$ good 6.73%
$r_{10}$ med $\rightarrow$ high 1 $\rightarrow$ 0 bad $\rightarrow$ good 4.36%
$r_{11}$ med $\rightarrow$ high 1 bad $\rightarrow$ good 3.96%
@@
$r_{19}$ mechan. med $\rightarrow$ high med $\rightarrow$ high 1 $\rightarrow$ 0 0 $\rightarrow$ 1 bad $\rightarrow$ good 1.51%
@@ 4 Discussion
Feature generation We used the base open-source LLama2 language model for a generation of text characteristics based on abstracts. These features are directly derived from the text, meaning, and context provided in the article abstracts. Many of these features would have to be assessed by a human evaluator, making the use of LLM a natural candidate for their extraction.
Predictive performance compared to existing representations First, we tested on two different datasets, which resulted in a marked increase in the accuracy over the baseline naive classifiers (50.2% vs 59.8% for CORD-19, 18% vs 37% for M17+). This showed that LLM-generated features are informative, but how do they stand in comparison with standard text representation? In the case of predicting human evaluation (M17+), the model trained on LLM-generated features surpassed the classical TF-IDF-weighted BoW approach in terms of accuracy and was tied on the F1 score and MAE.
@@
Analysis of importance of individual LLM features Interpretation of machine-learning models using SHAP indicated that research rigor has the highest importance in predicting the citation rate (CORD-19) as well as in predicting human evaluation (M17+). Interestingly, the effects of LLM-generated features were more pronounced in their relation to the human evaluation rather than citation rate. This is likely due to the human evaluation being performed by a jury of experts that take these characteristics of text into account, whereas previous bibliometric research [15] indicates that the number of citations is also influenced by the bibliometric features such as journal impact factor or the number of authors, which are features that we did not include into our analysis.
@@
Rule learning over LLM-based features In our previous research [4], which involved rule mining over the BoW text representation of a subset of CORD-19 dataset, we struggled with the problem of highly specialized rules referring to individual rule conditions. This was partly addressed by rule clustering, which grouped similar rules together, but this had considerable limitations. The use of more general LLM-based features is a leap forward, as it is possible to define features at an arbitrary-level of generality. For example, we had LLM generate a binary feature indicating, whether the article abstract refers to any experiments. In [4], we could infer the presence of experiments only from specific words, such as ”mouse” that would suggest that experiments using murine models on animals were conducted. Despite the presence of better features, the rule-based experiments in the present article exhibited some redundancy, but not on the level of rules. In the results, we often observed similar rules, which often different only in one or two conditions and covered the same or highly similar set of articles. To counter this problem, the concept of dominant action rules was implemented, which removed rules representing mere extensions of an existing rule without any improvement in uplift.
@@ 5 Related work
None of the earlier works performs a head-to-head comparison of the interpretability of LLM-based features and the bag of words (TF-IDF) features as we do. Crucially, none of the two previous works aimed to generate LLM-based features for white-box model learning. The key contribution of our work is that we demonstrated the effectiveness of the use of LLM-generated features in rule learning.
@@ 6 Conclusion
We provide an extensive evaluation of the LLM-based feature generation involving two datasets from the scientometric domain. The results overall show that the LLM-based features perform well and are interpretable. For example, already at the level of correlations with the target variable, LLM-based features such as rigour demonstrate an (expected) positive correlation with the number of citations a paper has received or how well it was evaluated by experts. The core of the evaluation is done by comparing the predictive performance of classifiers built with LLM-based features to those using other established feature representations and classifiers (TF-IDF, SciBERT embeddings, Autogluon). The results on the CORD-19 dataset show that LLM-based features result only in 3% lower accuracy compared to BoW (TF-IDF), while using much fewer features (62 compared to 1500). Additional reduction could be achievable through feature selection. On the M17 dataset, a classifier built with LLM-based features exceeds the accuracy of one built over TF-IDF. Interestingly, LLM features combined with BoW-weighted words perform on par with SciBERT embeddings. In addition to the quantitative results, we also demonstrated a practical use case, showing how LLM-based features can be used to build actionable rule-based predictions.
@@
Code and data is available from the authors at https://github.com/vojtech-balek/llm-features.
@@
V.B. data processing and analysis, L.S. action rule analysis, V.B., T.K. and L.S. writing, T.K. and V.B. methodology, V.S. scientometric expertise, M17+ dataset feature preparation, T.K. conceptualization, approval of the final manuscript, all authors.

### v2

@@ 3.3.6 Explanation of generated features
The explanation of features with the use of Shapley values was done on two pairs of classifiers. First, we compared a classifier trained on the results of a fully automated LLM-based feature generation with a classifier that was trained on user-selected features combined with LLM-based feature generation. This evaluation was done on the M17+ dataset. To complement this, we present the Shapley values for two additional datasets using the automated LLM feature extraction setup.
Results shown in Figure 1 indicate that for the M17+ dataset, the most important features from the user-selected ones were rigor, grammar, replicability, and novelty. For a classifier trained on fully automatic feature extraction, the most important features are language complexity, research discipline engineering, research impact high, and complex methodology. The comparison of both results shows a match as the “manual” grammar feature is closely related to “automatic” language complexity, rigor is related to the complexity of methodology, and novelty is related to high research impact.
Interestingly, both analyses using SHAP values support our initial hypothesis by showing that higher novelty is linked to better expert evaluations. This association is further validated by the statistical tests and effect sizes discussed in previous sections.
In Figure 2 (left), the analysis shows that for the Hate Speech dataset and specifically for predicting class ’1’ (the presence of hate speech), the feature Presence of Racial Slurs_Yes has the highest impact on the model’s output. Additionally, the SHAP values indicate that non-neutral sentiment and the use of emotive language are strongly associated with hate speech, as could be expected.
Figure 2 (right) shows that the most important LLM-generated feature is the LLM’s attempt to assess the target label for the Food Hazard dataset directly, as we can see from the feature hazard_type_biological. It can also be observed that the presence of Salmonella and the stated recall reason being contamination are intuitively understandable as important features, also for non-experts.
@@
Figure 1: SHAP plot for a classification model trained on LLM-feature only feature set for the M17+ dataset: automated LLM-based feature discovery combined with LLM-based feature generation (left) vs user-selected features combined with LLM-based feature generation (right). The X-axis represents the SHAP value, indicating the feature’s effect on the prediction—positive values increase the prediction, while negative values decrease it. Features are ranked by importance on the Y-axis. The color of the points corresponds to the feature values. For binary features like grammar or basic_medicine, blue indicates a value of 0, while red represents a value of 1. For categorical features such as rigor or novelty, blue denotes a low value, purple indicates a medium value, and red signifies a high value.
Figure 2: SHAP plot for classification models trained on LLM-feature only feature set for the Hate Speech dataset, target class 1 (hate speech present) (left) and the Food Hazard dataset target class 1 (biological food hazard) (right). For guidance on the general interpretation of the SHAP plot, refer to the caption of the previous figure.
3.3.7 User-perceived relevance of LLM-discovered features
To complement the quantitative evaluation of automatically discovered features by evaluating their predictive performance and match with domain-specific feature generation methods, we also conducted a small-scale study of human perception of the generated features.
To evaluate the appropriateness of automatically discovered features, we surveyed $41$ respondents. Participants self-identified as Business Professionals, Academic/Researchers, Undergraduate Students, or Graduate Students, and reported their machine learning expertise as None, Low (minimal understanding, no practical use), Moderate (understands core concepts with limited practice), or High (strong theoretical/practical experience), as shown in Figure 3(a).
Each participant was randomly presented a brief description of one of five distinct datasets and its corresponding prediction task, for which they were asked to rate the relevance of 20 automatically discovered features. Similarly to [15], we asked the participants to rate feature relevance on a 5‑point relevance scale (1 = “Not relevant” to 5 = “Relevant”).
The distribution of results, shown in Figure 3(b), is positively skewed toward higher values. Of the 100 features that were evaluated in total across the five datasets, only four features were considered as clearly not relevant (mean score between 1.0 and 2.0). In contrast, 27 features were evaluated as clearly relevant (mean score 4.0 and 5.0). Similarly, 50 features had a mean relevance score between 3.0 and 4.0, while 19 had a mean score between 2.0 and 3.0. The low number of features considered as not relevant suggests that LLM rarely outputs completely irrelevant or “hallucinated” features. With a substantial proportion of discovered features being judged as relevant by the evaluators, the results support the efficacy of our automatic generation pipeline.
(a) Survey participants ($N=41$) by role and machine learning knowledge.
(b) Histogram of mean relevance scores (1 = “Not relevant” to 5 = “Relevant”) for N=100 evaluated features pooled from 5 datasets.
Figure 3: Survey of user-perceived relevance of automatically discovered features. Each survey participant rated the relevance of 20 features that were automatically discovered for one of five datasets (CORD-19, M17+, BANKING77, Hate Speech and Food Hazard).
3.4 Case Study: Action Rule Mining for Article Improvement
We focus on applying action rules on scientometric datasets as these present the most promising opportunities for interventions proposed by action rules. For the analysis in this section, we used the CORD-19 and M17+ dataset versions with user-defined features.
@@ 3.4.1 M17+ Dataset - Bad to Good Evaluation Transition
$r_{9}$ low $\rightarrow$ high 0 $\rightarrow$ 1 bad $\rightarrow$ good 6.73%
$r_{10}$ low $\rightarrow$ high 1 $\rightarrow$ 0 bad $\rightarrow$ good 4.36%
$r_{11}$ low $\rightarrow$ high 1 bad $\rightarrow$ good 3.96%
@@
$r_{19}$ mechan. med $\rightarrow$ high 1 $\rightarrow$ 0 0 $\rightarrow$ 1 bad $\rightarrow$ good 1.51%
@@ 3.4.3 M17+ Dataset - Area-Specific Evaluation Improvement
Rules $r_{6}$ and $r_{7}$ stand out due to their ability to reclassify nearly 17% of articles in the entire dataset (16.95% uplift) from the undesired target class Evaluation = bad to the desired target class Evaluation = good. This highlights the substantial potential of these action rules to positively influence article evaluation outcomes on a large scale.
@@ 4 Discussion
Feature generation We used the base open-source Llama2 language model for the generation of text characteristics based on abstracts. These features are directly derived from the text, meaning, and context provided in the article abstracts. Many of these features would have to be assessed by a human evaluator, making the use of LLM a natural candidate for their extraction.
Predictive performance compared to existing representations First, we tested on two different datasets, which resulted in a marked increase in the accuracy over the baseline naive classifiers (50.2% vs 59.8% for CORD-19, 18% vs 37% for M17+). This showed that LLM-generated features are informative, but how do they stand in comparison with standard text representation? In the case of predicting human evaluation (M17+), the model trained on LLM-generated features surpassed the classical TF-IDF-weighted BoW approach in terms of accuracy and was tied for the F1 score and MAE.
@@
Analysis of importance of individual LLM features Interpretation of machine-learning models using SHAP indicated that research rigor has the highest importance in predicting the citation rate (CORD-19) as well as in predicting human evaluation (M17+). Interestingly, the effects of LLM-generated features were more pronounced in their relation to the human evaluation rather than citation rate. This is likely due to the human evaluation being performed by a jury of experts who take these characteristics of text into account, whereas previous bibliometric research [29] indicates that the number of citations is also influenced by the bibliometric features such as journal impact factor or the number of authors, which are features that we did not include in our analysis.
The quality of LLM-generated features on the scientometric datasets could be further improved if full texts were used instead of abstracts, although the magnitude of this improvement is hard to predict. For example, [12] observed only a slight improvement.
Accuracy-Interpretability tradeoff LLM-based feature generation is favourably positioned on the interpretability-accuracy spectrum. The interpretability is its core strength - the generated features have user-friendly and intuitively important features like Statistical analysis - None (important for the M17+ dataset predicting research impact) or Presence of Racial Slurs - Yes (important for the Hate speech detection dataset). These names can also be optionally described by the additional text, which could be very useful in the case of highly specialised features such as those present in the Food hazard dataset. This is an advantage even compared to features commonly present in off-the-shelf datasets used for machine learning, which often lack these qualities [13, 27]. In terms of accuracy, our results suggest that while classifier performance on LLM features may not reach the heights of deep learning methods that represent data in the latent space of embeddings, LLM features performed equally well or better against common interpretable baselines, such as TF-IDF.
It should be emphasized that LLM feature generation provides large opportunities for further improvements. For example, in our experiment focusing on the computational cost benchmark (Section 3.1.1), we used a chain of thought DeepSeek reasoning model. While it was out of the scope of this research to analyze its outputs in detail, an analysis of reasoning for excluding or including a specific feature could provide a viable alternative or complement to traditional approaches based on statistical feature selection, such as chi-square significance testing of forward feature selection. While not necessarily providing better results, the textual reasons will be better understandable for most users than numerical test results.
@@
Rule learning over LLM-based features In our previous research [5], which involved rule mining over the BoW text representation of a subset of CORD-19 dataset, we struggled with the problem of highly specialized rules referring to individual rule conditions. This was partly addressed by rule clustering, which grouped similar rules together, but this had considerable limitations. The use of more general LLM-based features is a leap forward, as it is possible to define features at an arbitrary level of generality. For example, we had LLM generate a binary feature indicating whether the article abstract refers to any experiments. In [5], we could infer the presence of experiments only from specific words, such as ”mouse” that would suggest that experiments using murine models on animals were conducted. Despite the presence of better features, the rule-based experiments in the present article exhibited some redundancy, but not on the level of rules. In the results, we often observed similar rules, which were often different only in one or two conditions and covered the same or highly similar set of articles. To counter this problem, the concept of dominant action rules was implemented, which removed rules representing mere extensions of an existing rule without any improvement in uplift.
@@ 5 Related work
[…]
Comparison to other LLM-based feature generation methods
@@
None of the earlier works performs a head-to-head comparison of the interpretability of LLM-based features and the bag of words (TF-IDF) features as we do. Crucially, neither of the two previous works aimed to generate LLM-based features for white-box model learning. The key contribution of our work is that we demonstrated the effectiveness of the use of LLM-generated features in rule learning.
Comparison to alternative non-LLM feature generation methods
The literature covers a large variety of feature extraction and generation methods. To scope the comparison, we focus on those that were used in conjunction with the datasets included in our study.
Our prior research, [5, 11], focused on building interpretable classifiers for the CORD-19 and M17+ datasets. These works utilized mainly rule-based methods and logistic regression; both would benefit from a less-dimensional and more interpretable feature set generated by LLMs.
@@
[…]
In contrast to most aforementioned approaches, the LLM-driven method described in this article generates interpretable features through prompt-based interaction without dataset-specific engineering.
@@ 6 Conclusion
We have demonstrated a novel and effective approach using LLMs to extract low-dimensional, interpretable features from text, overcoming key limitations of traditional representations for interpretable machine learning. Our findings across diverse datasets confirm that these features offer competitive predictive performance compared to standard baselines like TF-IDF and even SciBERT embeddings, while critically retaining semantic meaning. This was particularly evident in the successful generation of understandable and actionable rules, demonstrating the practical utility for tasks demanding transparency. Both fully automated and user-specified feature generation workflows proved viable, offering flexibility for different use cases. This LLM-driven feature engineering paradigm opens promising avenues for applying a wider range of interpretable ML techniques directly to text.
@@
Code and data are available at https://github.com/vojtech-balek/llm-features.
@@
V.B. data processing and analysis, L.S. action rule analysis and feature discovery pipeline, V.B., T.K., and L.S. writing, T.K., L.S., and V.B. methodology, V.S. scientometric expertise, M17+ dataset feature preparation, T.K. conceptualization, approval of the final manuscript, all authors.
@@
Declaration of generative AI and AI-assisted technologies in the writing process.
During the preparation of this work, the authors used Grammarly, Gemini, and ChatGPT in order to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed.
Appendix A LLM prompts
"system_message": "IMPORTANT: Return only a valid JSON object with no explanations, text, or markdown!!! Do not include any commentary or introductory text!!!",
"input_metadata": {
"dataset_name": "\$name",
"description": "\$description",
"target": "\$target",
"examples": "\$examples"
"task": {
"steps": [
"Analyze the provided metadata and examples to determine the domain and context of the dataset.",
"Identify the key characteristics of the dataset relevant to predicting the target variable.",
"List potential high-level categorical and numerical features based on domain knowledge inferred from the dataset description.",
"Extract additional potential features from dataset examples using syntactic and semantic patterns, ensuring at least 20 distinct features are generated.",
"If the text implies certain values that match the target, these values may also be extracted as features. In cases where the target has multiple values, each value can be independently derived from the text as a feature if it is contextually appropriate.",
"For text-based datasets, identify key phrases, structural components, and linguistic patterns that are relevant.",
"For numerical datasets, identify aggregation patterns, distributional characteristics, and possible transformations.",
"Group related features into meaningful categories where applicable.",
"If a feature has more than 15 unique categories, group less frequent categories into an Other class.",
"For each identified feature, provide a clear name, description, a complete list of possible values, and a specific LLM extraction query."
"constraints": [
"Ensure features are distinct and non-redundant.",
"Note that the target variable is not explicitly present in the input text.",
"Prioritize domain-specific insights over generic ones.",
"Ensure output is a structured, valid JSON format.",
"For categorical variables, list possible values with domain justification.",
"For numeric variables, provide possible transformations (e.g., log, mean differences).",
"The extraction queries must be specific and detailed to ensure high-quality feature generation.",
"Tailor extraction queries to the domain context of the dataset.",
"Generate a diverse set of features to maximize potential predictive power."
"output_format": {
"type": "json",
"structure": {
"features": [
"feature_name": "<Name of the categorical or numerical feature>",
"description": "<Short description of what the feature represents and how it relates to the datasets context>",
"possible_values": ["<Value 1>", "<Value 2>", "...", "<Value n>"],
"extraction_query": "Identify the ’<feature_name>’ based on the provided context. Options: ’<Value 1>’, ’<Value 2>’, ..., ’<Value n>’."
Figure 4: Full prompt for LLM-based feature discovery. Prompt can be parameterised by inserting specific values instead of placeholders, which are denoted with \$ symbol.
"input_text": "\$text",
"task": "Extract the following features as described below and return a valid JSON object.",
"constraints": [
"The output must be a valid JSON.",
"All answers must be simple and correspond to categorical values only."
"features": [
"feature_name": "product_name",
"description": "The name of the product being recalled, which can indicate the type of food and potential hazard.",
"extraction_query": "Identify the product_name based on the provided context. Options: Dole Fresh Blueberries’, XSell Mini Vegetable Samosas’, ..., Shikar brand’."
"feature_name": "hazard_type",
"description": "The type of hazard associated with the recalled product, which can indicate the nature of the risk to consumers.",
"extraction_query": "Identify the hazard_type based on the provided context. Options: biological’, allergens’, chemical’, foreign bodies’, fraud’, packaging defect’, food additives and flavourings’, other hazard’."
"feature_name": "recall_reason",
"description": "The specific reason for the recall, which provides insight into the nature of the defect or contamination.",
"extraction_query": "Identify the recall_reason based on the provided context. Options: Cyclospora contamination’, undeclared milk, soya & wheat’, ..., incorrect packaging’."
"feature_name": "distribution_area",
"description": "The geographical area where the recalled product was distributed, which can indicate the scope of the recall.",
"extraction_query": "Identify the distribution_area based on the provided context. Options: United States’, Canada’, ..., Regional’."
"feature_name": "company_name",
"description": "The name of the company responsible for the recalled product, which can indicate the manufacturer or distributor involved.",
"extraction_query": "Identify the company_name based on the provided context. Options: Dole Diversified North America, Inc.’, Heron Foods’, ..., Universal Clearance Company’."
"feature_name": "product_category",
"description": "The category of the recalled product, which can indicate the type of food or beverage involved.",
"extraction_query": "Identify the product_category based on the provided context. Options: Fruits’, Vegetables’, ..., Packaged Foods’."
"feature_name": "recall_date",
"description": "The date when the recall was announced, which can indicate the timeliness of the recall action.",
"extraction_query": "Identify the recall_date based on the provided context. Format: YYYY-$M$-DD’."
"feature_name": "batch_code",
"description": "The batch code of the recalled product, which can help identify the specific production lot affected.",
"extraction_query": "Identify the batch_code based on the provided context. Provide the alphanumeric code."
"feature_name": "best_before_date",
"description": "The best before date of the recalled product, which can indicate the shelf life and potential risk period.",
"extraction_query": "Identify the best_before_date based on the provided context. Format: YYYY-$M$-DD’."
"feature_name": "consumer_advice",
"description": "The advice given to consumers regarding the recalled product, which can indicate the recommended actions to mitigate risk.",
"extraction_query": "Identify the consumer_advice based on the provided context. Options: Do not consume’, Return to store’, Dispose of product’, Contact company’."
"output_format": {
"type": "json",
"structure": {
"features": [
"feature_name": "<Feature Name>",
"answer": "<Extracted Answer>"
Figure 5: Example of an automatically generated prompt used for LLM-based feature generation for a single instance from the Food Hazard dataset. Prompt can be parameterised by inserting specific text instead of the placeholder \$text.